OpenStudy (anonymous):

What are the values of a, b, and c in the quadratic formula given the equation 3x^2 - x =4 ? a = 3, b = 0, c = -4 a = 3, b = -1, c = 4 a = 3, b = 1, c = 4 a = 3, b = -1, c = -4

OpenStudy (anonymous):

@imqwerty

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (anonymous):

my answer is c. is it right?

imqwerty (imqwerty):

Standard form of a quadratic equation is - ax^2 + bx+c=0 So we have 3x^2-x=4 , subtract, 4 from both sides Uget- 3x^2 - x-4=0 Compare this with ax^2+bx+c =0 a=3, b=-1, c=-4

OpenStudy (anonymous):

oh ok. Thanks bro

imqwerty (imqwerty):

No prblm :D